OpenStudy (anonymous):

In an industrial process, the temperature of a fixed amount of gas in a rigid container drops from 50 degrees Celsius to 25 degrees Celsius. What will be the final pressure in the container if the initial pressure is 3.6 atmospheres?

OpenStudy (aaronq):

\[\frac{ n _{2} R _{2} T _{2} }{ P _{2} V _{2} }=\frac{ n _{1} R _{1} T _{1} }{ P _{1} V _{1} } \rightarrow \frac{ T_{2} }{P _{2} }=\frac{ T _{1} }{ P _{1} }\]

OpenStudy (aaronq):

since moles, volume and R are constant